OpenStudy (anonymous):

Your desk is 4.5 feet wide. Find the equivalent width in meters.

OpenStudy (phi):

to change units you need a conversion factor for example 3.28 feet per meter written \( 3.28 \frac{\text{ft}}{\text{m} } \) you will want the feet to divide out, and you want meters "up top" if the factor is upside down (like here where we have feet up top and meters below) flip the factor: \[ \frac{ 1}{3.28} \frac{\text{m}}{\text{ft} } \] now multiply \[ 4.5 \cancel{\text{ ft}} \cdot \frac{ 1}{3.28} \frac{\text{m}}{\cancel{\text{ft}} } \] notice that feet divided by feet cancel (treat them like they were numbers, and anything divided by itself is 1) now simplify \[ \frac{4.5}{3.28} \text{ m} \]